Antacids are drugs that are used to combat hyperacidity by neutralizing the excess acid (HCl) present in the stomach. Another popular component of antacid preparations is aluminum hydroxide, Al(OH)3.

3.If the neutralization reaction is only 94.3 percent efficient, how many grams of HCI can be actually neutralized by 1 tablespoon of preparation?
1
Expert's answer
2020-11-25T02:21:53-0500

Al(OH)3 + 3HCl = AlCl3 + 3H2O

If 1 tablespoon will be 1 g of mass, then

n(Al(OH)3) = 1g/78g/mol = 0.0128 mol

According to the equation

n(HCl) = 3*n(Al(OH)3) = 3*0.0128 = 0.0384 mol

m(HCl) = 0.0384mol*36.5g/mol = 1.4 g - can be actually neutralized by 1 g tablespoon of Al(OH)3